fax.cls Documentation

J.B. Rhebergen

July 20, 1995

1 Introduction

This fax class is written to work with the following other classes or packages:

artikel3.cls One of the dutch article styles.
lastpage.sty A small package that lets you reference the last page.

It can however still work without these files. If the artikel3.cls cannot be found the article.cls will be loaded instead. If available lastpage.sty will also be loaded. If you do not have lastpage.sty you can set the fax page numbers by

hand by using the \setcounter{faxpages}{..} command.

2 Other files

The following files should be prepared by the user:

from.fax containing your name etc.
note.fax containing a note which you would like to add.

If you are using the dutch class option you need:

van.fax containing your name etc. (in dutch)
nb.fax containing a note which you would like to add. (in dutch)

Of course this can easily be changed or added for any other language. Feel free to do so.

3 Options

The following class options are at your disposal: dutch Switches to dutch language and reads the files van.fax and nb.fax instead

     of the default from.fax and note.fax

note Adds a note to your fax header. This note should be specified in the file

note.fax or if you have used the dutch option nb.fax nosep Omits the faxtopset length variable which is currently set at 3ex This

white space separates your fax header from the rest of the text. The fax is typeset in \sf (sans serif) by default, because I felt that this would give the most legible results on a fax machine.

4 Commands

In the preamble of the fax document the following commands are available: \to Specifies what you want in the ‘to’ box. This should preferably be in the form

of a tabular environment. If omitted the box will stay empty.

     \to{\begin{tabular}[t]{rl}
       to:      &You\\
       address: &Your place\\
       tel:     &Your number\\
       fax:     &Whatever\\
       \end{tabular}}

\from Specifies what you want in the ‘from’ box. This should preferably be in the form of a tabular environment. If omitted the from.fax file will be used.

     \from{\begin{tabular}[t]{rl}
       from:    &Me\\
       address: &My place\\
       tel:     &My number\\
       fax:     &Whatever\\
       \end{tabular}}

If you want an empty ‘from’ box use \from{}. \mymsg Normally the fax header begins with the message: Tele-Fax Message.

If you have specified the dutch option the fax header will begin with: Fax Boodschap If you want to start with your own header (or logo even) you can use the \mymsg command like this:

            \mymsg{{\Large\bsf Watch out! here comes a FAX!}}

If you want absolutely no fax top message just specify \mymsg{} i.e. an empty command.

\bsf Switches to bold sans serif. i.e. {\bsf this is bold} \slsf Switches to slanted sans serif. i.e. {\slsf this is slant}

5 Suggestions

To make your faxes even more ‘beautiful’ you could use the wasysym package which provides a \phone command. This can be used to get a cute little telephone icon in your ‘to’ or ‘from’ box. Like this:

\from{\begin{tabular}[t]{rl}
  from:    &Me\\
  address: &My place\\
  \phone:  &My number\\
  fax:     &Whatever\\
  \end{tabular}}

Don’t forget to put \usepackage{wasysym} in your preamble though.

For my own convenience I have made a signature package. This provides a signature for my fax documents. Of course you need to get your signature scanned first and make the signature available in the form of a .pcx file or .msp file. If you want to know more about this compile and read the sigdoc.tex file.

6 Conversion

Once you’ve successfully used the fax.cls to make a fax document you have a choice.

• Print it, feed it to the fax machine and send it. (trivial)

• Convert it to a fax-modem compatible file and send it. These days lots of people have a fax-modem. To convert the .dvi file you generated, to some kind of fax-modem compatible file you need a conversion program. Let me describe how I go about sending a fax.

I am using emTEX for DOS to compile my .tex documents. The latest emTEX release, version [4a], has a fax.cnf file and fax_base.fli file which facilitate generating .dvi files in fax resolution.

I subsequently convert this fax .dvi file to a .pcx file, using the following batch file.

@echo off
dvidrv dvidot pcx @fax.cnf /fl=-1 %1 %1.p?? %2 %3 %4 %5 %6 %7 %8 %9

This will generate a number of .pcx files one for each page of the fax document.

The .pcx files are numbered myfax.p01, myfax.p02, myfax.p03 if your original

myfax.dvi has three pages.

The last step is to convert these .p?? files to one single .fax file. I have found the 2FAX program by Hans Harder very useful. However there was a bug in this program. Herre de Jonge the coauthor of this fax.cls and I have debugged the program and added some functionality. Actually, Herre did most of the work on debugging the 2FAX program while I did most of the work on the fax.cls. We’ve modified version 1.62 of 2FAX and called it version 1.62a. I will make this debugged version available for ftp as soon as possible. We’ve also added some functionality. You now can use wildcards at the 2FAX command line. Like this: 2fax indoc.p?? outdoc.fax /hr

At this stage you have a file which can be sent by the fax-modem. In the case of 2FAX this file is in the ZFAX format. I am using an unregistered version of BGFAX to send (or receive) faxes. BGFAX is written by B.J.Guillot and is a simple but solid command line driven fax program. It’s primarily intended for use with people who operate a bulletin board but who also want to send or receive faxes. There are various other programs for DOS that will also send ZFAX files. I will probably send the debugged version of 2FAX to B.J. Guillot’s bulletin board soon so people can download or ftp it from there. At the moment I still have to write a readme file and make some minor changes. If you’re in a hurry to get the debugged version of 2FAX, mail me and I’ll send it to you.

jbrhebergen@et.tudelft.nl
rheberg@morra.et.tudelft.nl